The mean free path of photon stars is is of the order of 0.2 centimetre. Show that the time taken for a Photon to reach the surface of a star of radius 4R is of the order of 1 million year.
1
Expert's answer
2020-04-21T10:59:58-0400

The distance d moved by a photon that random walks through N steps:


"4R=l\sqrt{N}"

We have then the following number of steps


"N=\frac{16R^2}{l^2}"

"T=\frac{16R^2}{l^2}t=\frac{16R^2}{lc}"

"T=\frac{16(695700)^2}{(0.2\cdot10^{-5})(299792)}=1.3\cdot10^{13}s=0.41\cdot10^{6}years"